you’re driving a car, traveling at a constant velocity of 25m/s along a straight road, when you see a child suddenly run into a road. it takes 0.45 s for you to react and apply the breaks. as a result the car slows with an acceleration of 8.5m/s^s in the direction opposite your motion and comes to a stop. what is the total displacement of the car before it stops?
1
Expert's answer
2020-10-05T10:53:30-0400
"S=s+d\\d=vt=(25)(0.45)=11.25\ m\\v^2=2as\\25^2=2(8.5)s\\s=36.76\ m\\S=11.25+36.76=48.0\ m"